A constant volume gas thermometer records a pressure of 250mmHg at 0°c and 290mmHg at 100°c. Calculate the room temperature when the gas pressure is 262mmHg
1
Expert's answer
2020-05-12T09:58:34-0400

Solution.

"p_1=250mmHg;"

"t_1=0^oC;"

"p_2=290mmHg;"

"t_2=100^oC;"

"V=const;"

"p=262mmHg;"

"t-?;"

Find the change in pressure when the temperature changes by 1oC:

"\dfrac{\Delta p}{\Delta t}=\dfrac{p_2-p_1}{t_2-t_1};"

"\dfrac{\Delta p}{\Delta t}=\dfrac{290mmHg-250mmHg}{100^oC-0^oC}=0.4\dfrac{mmHg}{^oC};"

"\dfrac{\Delta p}{\Delta t}=\dfrac{p-p_1}{t-t_1};\implies t-t_1=\dfrac{p-p_1}{\dfrac{\Delta p}{\Delta t}};"

"t-t_1=\dfrac{262mmHg-250mmHg}{0.4\dfrac{mmHg}{^oC}}=30^oC;"

"t=30^oC+0^oC=30^oC;"

Answer: "t=30^oC;"